OpenStudy (anonymous):

Find first partial sums S1 S2 S3 S4 and suggest a formula for Sn \[\frac{ 1 }{ 1*2 }+\frac{ 1 }{ 2*3 }+\frac{ 1 }{ 3*4 }+...+ \frac{ 1 }{ n(n+1) }+...\]

OpenStudy (anonymous):



OpenStudy (anonymous):

you welcome

OpenStudy (austinl):

The sum is, \[\frac{1}{n(n+1)}\] So the first term, you would simply plug 1 into it as "n".

OpenStudy (anonymous):

i suggest you actually add and see what you get as a partial sum

OpenStudy (austinl):

Well, it asks for S1-4... So you would need to solve out for those... right?

OpenStudy (anonymous):

\[S_1=\frac{1}{2}\] \[S_2=\frac{1}{2}+\frac{1}{6}=\frac{2}{3}\]

OpenStudy (austinl):

I will leave this to you, you seem like you have it well in hand.

OpenStudy (anonymous):

Wait, so i add them together?

OpenStudy (austinl):

For S2 you take what you got for n=1 and n=2, and add them together, yes.

OpenStudy (anonymous):

how do i find S4?

OpenStudy (austinl):

Well, S1 is n=1 S2 is n=1+n=2 S3 is n=2+n=3 S4 is n=3+n=4 Does this make sense?

OpenStudy (anonymous):

Why does n=1?

OpenStudy (austinl):

I assume this is what you are doing? \[\sum_{1}^{\infty} \frac{1}{n(n+1)}\]

OpenStudy (anonymous):

Oh! I get it! Thank you! :)